# Structure of the mean-field yrast spectrum of a two-component Bose gas in a ring: role of interaction asymmetry

### v1

@@ IV.2 $\kappa>0$
Figure 10: Critical curves for plane-wave yrast states $(\phi_{0},\phi_{k})$ with $k=2,3$ and the corresponding phase diagram for $\kappa=1$.
@@
where $\lfloor q/2\rfloor$ denotes the integer part (floor) of $q/2$. Once $k$ lies within this range, then the previous argument implies that $(\phi_{0},\phi_{k})$ will eventually become an yrast state for sufficiently large $\gamma$. Consequently, for a fixed $x_{B}=p/q$, the plane-wave states $(\phi_{0},\phi_{k})$ emerge as yrast states one by one with increasing $\gamma$, in order of increasing $k$, up to $k=\lfloor q/2\rfloor$. Equation  30 also means that the plane-wave state $(\phi_{0},\phi_{k})$ can never become an yrast state for those concentrations $x_{B}=p/q$ satisfying $\lfloor q/2\rfloor<k$, regardless of how large $\gamma$ is. This implies, for example, that $(\phi_{0},\phi_{2})$ can never become an yrast state at $x_{B}=1/2$ or $x_{B}=1/3$, which is precisely why these values appear as asymptotes of the critical curve $x_{B}(\gamma,k=2)$. Likewise, $(\phi_{0},\phi_{3})$ can never become an yrast state for $x_{B}=1/2,1/3,1/4,1/5,$ and $2/5$. This conclusion is confirmed by our numerical determination of the critical curve $x_{B}(\gamma,k=3)$, shown along with a partial phase diagram in Fig. 10, where these values indeed serve as asymptotes. The asymptotes of the critical curves for higher $k$ values can be deduced similarly.
Importantly, the critical curves determined here deviate substantially from those obtained perturbatively in Ref. [Wu2015]. This indicates that, for $\kappa>0$, plane-wave yrast states do not emerge through a continuous transformation from soliton states, as is the case for $\kappa\leq 0$. To substantiate this observation, we examine the evolution of the condensate wave functions at $l=2x_{B}$ for $x_{B}=0.2$ and $\kappa=1$ across the critical value $\gamma_{c}\simeq 5.2$, as indicated in Fig. 10. In Fig. 11, we plot the condensate wave functions at $\gamma=5$ and $\gamma=5.5$, corresponding to values just below and above the critical point. We observe that the condensate wave function at $\gamma=5$ is a soliton state that does not continuously evolve into the plane-wave state observed at $\gamma=5.5$. This clearly demonstrates that the emergence of the plane-wave yrast state occurs via a branch crossing rather than through a continuous transformation.
Figure 11: Plots of the amplitude as a function of the phase in the polar coordinate, displayed in the bottom panel.

### v2

@@ IV.2 $\kappa>0$
Figure 10: The critical curves for plane-wave yrast states $(\phi_{0},\phi_{k})$ with $k=2,3,4$ and the corresponding phase diagram for $\kappa=1$.
Figure 11: The upper and middle panels, respectively, show the amplitude and phase of the condensate wave function at $l=2x_{B}$ with $x_{B}=0.2$ and $\kappa=1$ for $\gamma=3,4,5$ and $5.5$ (see the dots on the dashed line with arrow in Fig. 10). The bottom panel plots the amplitude of the condensate wave function against the phase in the polar coordinate.
@@
where $\lfloor q/2\rfloor$ denotes the integer part (floor) of $q/2$. Once $k$ lies within this range, then the previous argument implies that $(\phi_{0},\phi_{k})$ will eventually become an yrast state for sufficiently large $\gamma$. Consequently, for a fixed $x_{B}=p/q$, the plane-wave states $(\phi_{0},\phi_{k})$ emerge as yrast states one by one with increasing $\gamma$, in order of increasing $k$, up to $k=\lfloor q/2\rfloor$. Equation  30 also means that the plane-wave state $(\phi_{0},\phi_{k})$ can never become an yrast state for those concentrations $x_{B}=p/q$ satisfying $\lfloor q/2\rfloor<k$, regardless of how large $\gamma$ is. This implies, for example, that $(\phi_{0},\phi_{2})$ can never become an yrast state at $x_{B}=1/2$ or $x_{B}=1/3$, which is precisely why these values appear as asymptotes of the critical curve $x_{B}(\gamma,k=2)$. Likewise, $(\phi_{0},\phi_{3})$ can never become an yrast state for $x_{B}=1/2,1/3,1/4,1/5,$ and $2/5$. This conclusion is confirmed by our numerical determination of the critical curve $x_{B}(\gamma,k=3)$, shown along with a partial phase diagram in Fig. 10, where these values indeed serve as asymptotes. The asymptotes of the critical curves for higher $k$ values can be deduced similarly. Once again, as in the case of $\kappa<0$, the phase diagram shown in Fig. 10 provides a clear explanation for the nonanalytic structure of the yrast spectrum displayed in Fig. 5.
Figure 12: The two lowest spectral branches in the vicinity of $l=2x_{B}$, distinguished by the different winding numbers of the condensate wave functions, are shown for $\gamma=5$ (left) and $\gamma=5.5$ (right). At $\gamma=5$, the branch containing a soliton state at $l=2x_{B}$ constitutes the yrast spectrum, as it lies below the branch containing the plane-wave state $(\phi_{0},\phi_{2})$ (marked by the red square). As $\gamma$ increases, the plane-wave branch shifts downward relative to the soliton branch, and the plane-wave state $(\phi_{0},\phi_{2})$ becomes degenerate with the soliton state in the other branch at the critical value $\gamma_{c}\simeq 5.2$. The crossing of the two branches is clearly visible in the right panel at $\gamma=5.5$.
Importantly, the critical curves determined here deviate substantially from those obtained perturbatively in Ref. [Wu2015]. This indicates that, for $\kappa>0$, plane-wave yrast states do not emerge through a continuous transformation from soliton states, as is the case for $\kappa\leq 0$. To substantiate this observation, we examine the evolution of the condensate wave functions at $l=2x_{B}$ for $x_{B}=0.2$ and $\kappa=1$ across the critical value $\gamma_{c}\simeq 5.2$, as indicated in Fig. 10. In Fig. 11, we plot the condensate wave functions at $\gamma=3,4,5$ and $\gamma=5.5$, corresponding to values just below and above the critical point. We observe that the condensate wave functions at $\gamma=3,4$ and $5$ are soliton states that do not continuously evolve into the plane-wave state $(\phi_{0},\phi_{2})$ found at $\gamma=5.5$. This indicates that the emergence of the plane-wave yrast state occurs via a branch crossing rather than through a continuous transformation. This process is illustrated explicitly in Fig. 12, where two lowest branches of spectrum —one containing the plane-wave state $(\phi_{0},\phi_{2})$ at $l=2x_{B}$ and the other a soliton state—cross as $\gamma$ is varied between $5$ and $5.5$.